OpenStudy (anonymous):

4x-9y=-36 into slope and intercept form

OpenStudy (domebotnos):

4x - 9y = -36 -4x -4x -9y = -4x - 36 ____ ____ ____ -9 -9 -9 y = 4/9 + 4

OpenStudy (domebotnos):

Medal, plz?

geerky42 (geerky42):

No.